Best Buy Inc. has identified an investment project with the following cash flows. If the discount rate is 6%, what is the future value of these cash flows in year 4? If the discount rate is 3%, what is the future value in year 4? Show your work

Year 1. 728

Year 2. 1,032

Year 3. 1,346

Year 4. 1,674
Answer #1

Discount rate 6%:

Future value of year 1 cash flow = 728 (1 + 6%)3 = 867.0596

Future value of year 2 cash flow = 1,032 (1 + 6%)2 = 1,159.5552

Future value of year 3 cash flow = 1,346 (1 + 6%)1 = 1,426.76

Future value of year 4 cash flow = 1,674 (1 + 6%)0 = 1,674

Future value = 867.0596 + 1,159.5552 + 1,426.76 + 1,674

Future value = $5,127.37

Discount rate 3%:

Future value of year 1 cash flow = 728 (1 + 3%)3 = 795.505256

Future value of year 2 cash flow = 1,032 (1 + 3%)2 = 1,094.8488

Future value of year 3 cash flow = 1,346 (1 + 3%)1 = 1,386.38

Future value of year 4 cash flow = 1,674 (1 + 3%)0 = 1,674

Future value = 795.505256 + 1,094.8488 + 1,386.38 + 1,674

Future value = $4,950.73